OpenStudy (anonymous):

what was political change in the north during the 1800's

OpenStudy (anonymous):

population ****

sammixboo (sammixboo):

What north are you talking about?

sammixboo (sammixboo):

North can be anywhere North

TheSmartOne (thesmartone):

\[\begin{array}{lr} \,\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \overset{\bcancel{\;\,}}\rhd \overset{\cancel{\;\,}}\lhd\\ \underset{\widehat\qquad}\blacktriangle\\ \Large \color{#CD5555}{\overline{\quad}} \end{matrix} \right)}}\normalsize \\ \begin{array}{lcr} /&\color{blue}\bowtie&\backslash\\ |&\scriptsize^\circ&|\\ |&\scriptsize^\circ&|\\ \end{array} & \begin{array}{l} \large \text{You have posted \in the}\\ \large\text{wong area. Please read }\\ \large\text{the Code of Conduct.} \end{array} \end{array}\]

TheSmartOne (thesmartone):

@sammixboo

TheSmartOne (thesmartone):

I didn't make it... e.mccormick made it xD

OpenStudy (anonymous):

@jim_thompson5910